# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

### v2

@@ 2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is. Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
The used voting strategy combines the constructed lists of text chunks from individual embedding spaces into a single list of top $k$ chunks. The strategy is very simple (the corresponding Algorithm 2 is in the Appendix). Each text chunk from a list $i$ of the vector space $i$ has a certain position on this list, we denote this position with $p$. We obtain a weight for this chunk as $s_{i}\cdot 2^{-p}$; $s_{i}$ is the previously defined importance score of the space $i$. Multiplying $s_{i}$ with $2^{-p}$ exponentially lowers the significance of less relevant text chunks. Finally, all chunks from all lists are sorted using their weights and the top $k$ chunks form the final list.
2.3.1 Integration with Data Stores & Other Types of Models
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG. Similarly, MRAG does not depend specifically on the embedding form, as long as it is based on a model that harnesses multi-head attention any such approach that results in a valid embedding can be used. As such, it could also be used with models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024a).
3 Multi-Aspect Datasets, Queries, and Metrics
To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving different documents from the multi-aspect dataset, and (3) metrics that asses how well a given RAG scheme retrieves such multi-aspect data. We now describe these three elements. In Section 4, we also discuss real-world data and queries used to evaluate MRAG.
3.1 Multi-Aspect Datasets
We first select conceptually different categories of documents for a synthetic dataset. Here, we harness publicly available Wikipedia articles. In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes. We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants. We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks. Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.). Examples of accident causes are natural disasters, human mistakes, or lack of proper training. We fully release these datasets to propel RAG research. Details on all three datasets can be found in the Appendix B.2. In our evaluation, we use a total of 13,750 documents.
3.2 Multi-Aspect Query Generation
We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo). We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
3.3 Metrics
We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. While less desired, it still increases chances for a more accurate LLM answer following the retrieval. For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$. By varying $w$, the user can adjust the importance of exact document matches and category matches. An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.
4 Evaluation
We now illustrate the advantages of MRAG over the state of the art.
Comparison Baselines We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024). The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG. Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.3).
We use queries and metrics introduced in Section 3. We use the weighted retrieval success ratio with 2:1 weighting, which considers category matches as relevant but prioritizes the exact document matches. Figure 3 shows an example query and metrics usage. Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth. We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches). Among these $k$ documents, we search for matches with the ground truth.
Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects. The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Section 3). A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
4.1 Superior Performance for Multi-Aspect Queries
We start from the query example in Figure 3 and show first the absolute retrieval performance of MRAG over Standard RAG in Figure 4. We fix the number of aspects present in the queries to 10, and vary the total number of retrieved documents from 10 to 30. MRAG consistently outperforms Standard RAG ($>10\%$ increase in the retrieval success ratio on average for exact document matches). Moreover, the retrieval performance increase is even more significant on category matches ($>25\%$ increase in the retrieval success ratio on average). The performance increase is further detailed in the histograms on the right side. Here, for a specific number of documents fetched, MRAG’s histogram indicates a better distribution of retrieval success ratios (across all 25 queries).
Next, Figure 5 shows the relative weighted performance improvement of MRAG with respect to Standard RAG as we vary the number of aspects present in the queries. We show data for two different embedding models (SFR and e5). MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the number of aspects present in the replies, for both models.
4.2 Comparable Performance for Single-Aspect Queries
We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected. Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
[…]
Documents Fetched Multi-Aspect Dataset Legal Dataset Accidents Dataset
SFR e5 SFR SFR
1 24/25 25/25 24/25 25/25 24/25 24/25 25/25 25/25
2 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
3 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
4.3 Further Improvements with Additional Tokens
We now show that MRAG can be seamlessly integrated with other RAG approaches: We combine MRAG with Fusion RAG, representing RAG schemes that use an LLM (additional token cost) for more accurate retrieval. Fusion RAG uses an LLM to create a fixed number of questions about the RAG query. Each question is separately applied through an embedding model using Standard RAG. We apply MRAG’s approach to each of these questions and denote the combined scheme as Fusion MRAG. Red plots of Figure 6 show that both Fusion RAG and Fusion MRAG perform better than Standard RAG, on average gaining 10 to 30% in accuracy. Fusion MRAG performs consistently better than pure Fusion RAG, indicating that these optimizations can be combined together. However, both Fusion strategies introduce a greater variance than MRAG and additional costs in terms of compute, latency, and tokens.
4.4 Benefits from Multi-Head Attention Solely
We also compare MRAG to the Split RAG baseline in Figure 6. The blue plots show the relative weighted performance of MRAG and Split RAG over Standard RAG. MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
4.5 Real-World Workloads
To further illustrate advantages of MRAG, we also consider two real-word use cases from in-house industry data analytics projects, namely, the synthesis of legal documents and the analysis of causes of chemical plant accidents. The results are in Figure 7. In the former (the left side), the task is to create a document based on user requirements that may be related to different aspects, for example to the law being considered (e.g., the British or the US one), the subject (e.g., energetic or civil), the style of the document (e.g., aggressive or mild), etc.. This task is executed with RAG that can fetch documents from a database. In the latter (the right side), the task is to discover a cause of an accident. Here, one also wants to retrieve documents from a database that should be used in the LLM context to facilitate discovering the cause of the accident. The causes are grouped in categories such as utility impact due to severe weather, lack of preparedness and planning, incorrect installation of equipment, lack of maintenance, etc.. Similarly to the previous analyses, we measure the retrieval success ratio over corresponding databases. MRAG offers advantages over other schemes.
4.6 Additional Analyses
We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one). Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3). We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
The complexity of the importance score calculation (Algorithm 1) is $\mathcal{O}(n^{2})$ where $n$ is the number of the embedded documents; it is dominated by calculating the pair-wise cosine similarity and the calculation of the norm. Please note that this step needs to be done only once for each dataset and is not a bottleneck.
Finally, in addition to the performance evaluation, we also investigated the attention heads of the SFR-Embedding-Mistral model as well as LLaMA-2 7B model (model not fine-tuned for text-embedding tasks). This analysis is presented in Appendix C.
5 Related Work
Our work touches on many areas which we now briefly discuss.
Many RAG schemes appeared recently (Gao et al., 2024), using the output of the last decoder layer for embedding generation. In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries. As such, it can be combined with other schemes to further improve RAG pipelines.
Retrieval is sometimes enhanced by a cross-encoder reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2021; Gao et al., 2021; MacAvaney et al., 2019). In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models. In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality. A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships (Jiang et al., 2024; Delile et al., 2024; Hussien et al., 2024; Bui et al., 2024; Xu et al., 2024a). RAPTOR (Sarthi et al., 2024) generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text. Graph RAG (Edge et al., 2024) creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction. All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text. Usually, they need a sophisticated preprocessing phase to prepare such structures. MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.
6 Conclusion
Retrieval Augmented Generation (RAG) is pivotal for democratizing access to accurate and relevant outputs from large language models (LLMs). Enhancing the precision and relevance of these outputs is a critical goal, especially given the challenges posed by queries requiring the retrieval of multiple documents with significantly different contents. These complex queries are common across various domains, but existing RAG solutions struggle because the embeddings of the necessary documents can be far apart in the embedding space, complicating their retrieval.
To address this gap, we introduced Multi-Head RAG (MRAG), a novel scheme that leverages the activations from the multi-head attention layer of decoder models instead of the traditional feed-forward layer. This approach is grounded in the insight that different attention heads can capture distinct aspects of the data. By using these diverse activations, MRAG creates embeddings that better represent the multifaceted nature of data items and queries, thus enhancing the retrieval accuracy for complex, multi-aspect queries. The simplicity and versatility of this idea allow it to be seamlessly integrated into any modern RAG pipeline or data analytics framework.
Our comprehensive evaluation methodology, including specific metrics, synthetic datasets, and real-world use cases, demonstrates MRAG’s effectiveness. The results indicate a significant improvement in the relevance of retrieved documents, with up to 20% better performance compared to modern RAG baselines. This validates MRAG’s potential to handle the intricacies of multi-aspect queries effectively.
Moreover, MRAG proves to be both cost-effective and energy-efficient. It does not require additional LLM queries, multiple model instances, increased storage, or multiple inference passes over the embedding model. This efficiency, combined with the enhanced retrieval accuracy, positions MRAG as a valuable advancement in the field of LLMs and RAG systems. By addressing the challenges of multi-aspectuality in queries, MRAG paves the way for more reliable and accurate LLM applications across diverse industries.
Acknowledgements
We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.
Appendix
Appendix A Model Design: Additional Details
A.1 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
@@
     find best matching $k$ text chunks
     for each chunk $d_{i,p}$ with index $p$ in top $k$ do
@@
  sort $l$ using weights $w_{i,p}$; return top $k$ elems
Appendix B Evaluation Methodology: Additional Details
B.1 Compute Resources
B.2 Dataset Details
dataset #categories #topics #documents total #documents
Wikipedia 25 50 documents per category 1250
Legal Documents 25 25 per category 10 per topic 6250
Accident Reports 25 25 per category 10 per topic 6250
B.3 Prompt Template for the Synthetic Dataset Generation
Table 3: Prompt template for query generation.
Please create a story about the attached <number of articles> articles on the topics <list of titles>.
@@ B.3 Prompt Template for the Synthetic Dataset Generation
Appendix C Attention Head Analysis
@@
For both models, we found out that the attention patterns vary significantly between the different attention heads. Still, we encountered two distinct patterns. First, the diagonal lines in Figures 9(a) and 9(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x. We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other. Second, horizontal and vertical lines in Figures 9(c) and 9(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence. An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
@@
We also detail attention scores (after applying softmax) of selected heads in Figures 10 and 11, when the model is processing the last token of its input. We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads. An interesting pattern we encountered was that for the SFR-Embedding-Mistral model (see Figure 11), all heads’ attention spiked significantly on the first line-break in the input sequence - either positively or negatively. We conjecture that this is a consequence of how the embedding model was fine-tuned and its intended usage pattern: embedding queries are usually prepended with a retrieval instruction, which is terminated by a line-break. The model likely learnt to summarise the necessary information about this instruction inside the terminating line-break.

### v5

@@ C.1 Ranking Strategy Details
C.2 Ranking Strategy Details
The voting strategy used by MRAG in its reranker is pictured in Algorithm 2.
@@ C.2 Ranking Strategy Details
  find best matching $c$ text chunks
  for each chunk $d_{i,p}$ with index $p$ in top $c$ do
@@
 sort $l$ using weights $w_{i,p}$
 return top $k$ elements of $l$
C.3 Integration with Data Stores
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
Appendix D Specification of Prompts
D.1 Prompt Template for Reader
Prompt Template for Reader You are presented with a series of articles, each potentially addressing different aspects of a topic.
1. <article title>
[<metadata>]
<body>
2. <article title>
[<metadata>]
<body>
Task: Carefully analyze the articles. When formulating your response to the question below, identify the relevant aspects or claims made within each article. Construct your answer by comparing, contrasting, or synthesizing these points in a coherent and logically structured manner. Your response should be supported by specific references to the content of the articles. Where applicable, acknowledge differences in perspectives, data, or assumptions across the sources. Aim for clarity, precision, and concise reasoning grounded in the evidence provided.
Please answer the following question: <query text>
D.2 Prompt Template for the Synthetic Dataset Generation
Prompt Template for Query Generation Please create a story about the attached <number of articles> articles on the topics <list of titles>.
@@ D.2 Prompt Template for the Synthetic Dataset Generation
Appendix E Multi-Aspectuality with Attention Heads without Additional Training
In MRAG, we extract embeddings from the hidden representations immediately after the attention block in the last decoder layer, avoiding any fine-tuning. This decision is based on an existing hypothesis that attention heads in Transformer models naturally differentiate during training, each attending to distinct aspects of the input data distribution.
E.1 Literature Survey
This hypothesis has been substantiated across various Transformer families. Wang et al. (2025a) introduce the Local Learning Coefficient (LLC), a measure of training dynamics at the head level. They show that attention heads begin with similar behavior but quickly diverge into functionally distinct clusters, each specializing in different patterns, ranging from local structure to multigram token groups. Olsson et al. (2022) identified “induction heads” in GPT-style models: specific heads that attend to earlier repeated sequences (e.g., in patterns like “X … Y … X”), enabling the model to learn simple in-context repetition without additional supervision. Further studies observed heads that consistently attend to names, suppress repetition, or shift attention predictably to structurally aligned tokens (McDougall et al., 2024; Gould et al., 2024).
Research in BERT-style models further supports these trends. Clark et al. (2019) demonstrate that different heads attend to direct objects, nominal modifiers, or punctuation tokens. Kovaleva et al. (2019) identify broad attention patterns like “vertical” heads (focusing on special tokens) and “diagonal” heads (attending to adjacent words). Htut et al. (2019) show that many heads correlate with syntactic dependency arcs.
Notably, while many heads specialize in meaningful ways, others appear to contribute little to model performance. Voita et al. (2019) and Michel et al. (2019) show that a large fraction of heads can be pruned without substantial performance loss, suggesting that specialization tends to concentrate in a smaller subset of effective heads.
E.2 Analysis of Multi-Head Patterns
@@
For both models, we found out that the attention patterns vary significantly between the different attention heads. Still, we encountered two distinct patterns. First, the diagonal lines in Figures 7(a) and 7(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x. We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other. Second, horizontal and vertical lines in Figures 7(c) and 7(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence. An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
@@
We also detail attention scores (after applying softmax) of selected heads in Figures 8 and 9, when the model is processing the last token of its input. We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads. An interesting pattern we encountered was that for the SFR-Embedding-Mistral model (see Figure 9), all heads’ attention spiked significantly on the first line-break in the input sequence - either positively or negatively. We conjecture that this is a consequence of how the embedding model was fine-tuned and its intended usage pattern: embedding queries are usually prepended with a retrieval instruction, which is terminated by a line-break. The model likely learnt to summarise the necessary information about this instruction inside the terminating line-break.
@@
Appendix F Complexity Analysis: Additional Details
We now discuss additional details for our complexity analysis of the RAG schemes.
In Poly-encoders (Humeau et al., 2020) retrieval, we use $s(n)$ to represent additional attention evaluations needed to prepare the $m$ code embeddings and their final context embeddings with all $n$ documents.
In Lewis et al. (2020) retrieval, the work and depth are increased by the additional generator model evaluation.
In ColBERT (Khattab & Zaharia, 2020) retrieval, the work is increased linearly and depth logarithmically with the average query and document size due to the maxsim evaluations. For storage, the overhead scales linearly with the average size of the document, due to per-token embeddings.
In EMDR (Singh et al., 2021) retrieval, the work is increased $k+1$ times due to T5 encoders run on the top $k$ documents. This corresponds to a $D_{m}$ depth increase as the evaluations can be parallelized.
In Self-RAG (Asai et al., 2024), once the model decides to conduct retrieval using the retrieval token, it embeds the query, runs nearest $k$ search, and evaluates the model twice for each best-$k$ document to predict the issup, isrel, isuse labels. The depth increases only by $D_{m}$ as the evaluations can be parallelized.
In Chain-of-Note (Yu et al., 2024b), during retrieval the $k$ best documents are fetched and then each is summarized sequentially by generating notes, resulting in a $ks(n)$ work and depth increase, with $s(n)$ representing the cost to generate an average note.
In RAPTOR (Sarthi et al., 2024), preprocessing involves creating a document tree, with leaves representing documents, and parents containing summaries of children, a cost we represent with $s(n)$ both for work and storage. Retrieval involves traversing the tree, which we encapsulate in a different $s(n)$.
RAGraph (Jiang et al., 2024) preprocessing involves embedding the documents by a graph model and creating the toy graphs, which we account for using $W_{e}$ and $D_{e}$. The toy graphs also create additional space requirements, which we account for using $s(n)$. For retrieval, additional key information such as environment or position-aware codes is used, represented by another $s(n)$ in our notation.
In RQ-RAG (Chan et al., 2024), the retrieval may be decomposed into multiple separate queries that are generated by the model and then evaluated in sequence using standard techniques, resulting in increases in work and depth that we denote using $s(n)$.
In ActiveRAG (Xu et al., 2024), retrieval is conducted using standard techniques, followed by three separate model evaluations: the Knowledge Assimilation (KA), Self-Inquiry (Q), and Thought Accommodation (TA) agents. As TA is similar to the work of an LLM answering the query, we omit its cost, like in all other frameworks, resulting in triple the work. Q and KA can be evaluated in parallel, increasing the depth only twice.
HiQA (Chen et al., 2024d) combines multiple retrieval strategies, using vector similarity matching, elastic search with BM25, and keyword matching. We represent these with $s(n)$ for work. Note that as these are independent, the depth is not increased. For preprocessing, HiQA uses a Hierarchical Contextual Augmentor, which creates a data hierarchy and introduces an overhead we denote by $s(n)$. HiQA also stores more information alongside vectors, such as keywords, increasing requirements by $s(n)$.
GraphRAG (Edge et al., 2025) creates a knowledge graph during preprocessing, which we denote with $W_{e}$ and $D_{e}$. It then extracts communities and their summaries, which we account for using $s(n)$. These are stored and require additional space also denoted by a different $s(n)$. During retrieval, GraphRAG uses an LLM to rank all communities in parallel by how useful they are which we estimate using $s(n)$.
In Fusion RAG (Rackauckas, 2024) retrieval, $k$ queries are generated and for each a standard RAG is evaluated, together with a reranking achieved by another model evaluation, which we estimate as $s(n)$.
In Meta-chunking (Zhao et al., 2025b), the documents are preprocessed by splitting them into chunks based on perplexity or margin sampling. As the decision whether to split or combine sentences in a document is based on an LLM, the preprocessing requires $l_{d}$ evaluations with some $s(n)$ postprocessing. The storage requirements are also increased by a different $s(n)$ as documents are stored not as a single vector but multiple vectors based on chunks. For the same reason, retrieval requires more work, as the number of vectors is increased by $s(n)$.
In MoC (Zhao et al., 2025a), chunks are created based on different granularities. Similarly to Meta-chunking, this means a larger number of vectors resulting in increased storage requirements, and retrieval work by $s(n)$. As MoC also includes routing and meta-chunkers, the preprocessing cost is increased by $s(n)$.
In Parametric RAG (Su et al., 2025), documents are represented as deltas of parameters that can be applied to the model. During preprocessing, the model is fine-tuned on a given document, resulting in a considerable increase of $s(n)$ in work and depth. As parameters are considerably larger than the hidden dimension, storage is also increased by a different $s(n)$. During retrieval, standard RAG fetches the documents, and the original model needs to be updated with their parameters, increasing work and depth by $s(n)$.
SuperRAG (Yang et al., 2025) first embeds the documents in a knowledge graph, which we represent by $W_{e}$ and $D_{e}$, and then uses this knowledge graph in retrieval to index it using $W_{i}$ and $D_{i}$. These also include any reranking SuperRAG might do.
HiRAG (Huang et al., 2025a) creates a hierarchical knowledge graph used for indexing in retrieval. Similarly, to SuperRAG, this increases the preprocessing and retrieval costs. We include in these the additional description and report generation that HiRAG conducts.
Appendix G Full Specification of Benchmarking Multi-Aspectuality
We provide more details on how to benchmark multi-aspectuality in RAG. Figure 10 shows an example query and metrics usage. Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth. We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches). Among these $k$ documents, we search for matches with the ground truth.
G.1 Multi-Aspect Datasets
We first select conceptually different categories of documents for a synthetic dataset. Here, we harness publicly available Wikipedia articles. In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes. We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants. We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks. Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.). Examples of accident causes are natural disasters, human mistakes, or lack of proper training. We fully release these datasets to propel RAG research. Details on all three datasets can be found in the Appendix H.2. In our evaluation, we use a total of 16,500 documents.
G.2 Multi-Aspect Query Generation
We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-4o). We construct 160 queries with 1, 2, 3, 4, 5, 6, 10, 15, 20 and 25 aspects (1600 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 10.
G.3 Metrics
We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used Reranker scheme $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. While less desired, it still increases chances for a more accurate LLM answer following the retrieval. For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$.
An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 10.
Appendix H Evaluation Setup: Additional Details
H.1 Compute Resources
H.2 Dataset Details
dataset #categories #topics #documents total #documents
Wikipedia 80 50 documents per category 4000
Legal Documents 25 25 per category 10 per topic 6250
Accident Reports 25 25 per category 10 per topic 6250
Appendix I Evaluation: Additional Results
We provide additional empirical evaluation results.
I.1 Harnessing Different Decoder Blocks
We analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one). Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
I.2 Analyzing Different Voting Strategies
We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 11 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3).
We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 11 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
I.3 Analyzing Preprocessing Overhead
One-time head importance scoring in MRAG introduces minimal preprocessing overhead on top of the standard embedding scheme. The scoring consists of computing: (i) average L2 norms per embedding space, and (ii) average pairwise cosine distances among embeddings within each space. To assess practical overhead, we analyze six datasets using the SFR-Embedding-Model. The additional time required to compute importance scores is measured as a percentage of the original embedding time. For example, on SciFact, the overhead was just 2.7%, and on NFCorpus, only 1.75%. Even for moderate-scale corpora such as ArguAna (310 seconds total encoding time), the overhead remained under 5%. These results assume full pairwise distance computation across all chunks; in practice, the harnessed sampling makes them even lower.